OpenStudy (anonymous):

A player hits a foul ball with an initial vertical velocity of 70 ft/s and an initial height of 5 feet. The maximum height reached by the ball is _____ feet. (Round to the nearest foot and enter only the number.)

OpenStudy (anonymous):

was busy somewhere else. using third equation of motion \[\Large 2g(h-h_{0})=v_{f}^2-v_{i}^2\] here g=-32.2 ft/sec^2 vf=0 h0=5ft so \[\Large 2*(-32.2)(h-5)=0-(70)^2\] \[\Large h-5=\frac{70^2}{2*32.2}\] can you solve for h ??

OpenStudy (anonymous):

@AudrianaS

OpenStudy (anonymous):

does h=81.09?

OpenStudy (anonymous):

yes .you got it.

OpenStudy (anonymous):

okay thank you so much!!

OpenStudy (anonymous):

you are welcome:)